The diagonal of the floor of a square room is 6 meters.
What is the length of a side of the room? Round to the nearest tenth.

Answer by checkley79(3341)

YOU HAVE A RIGHT TRIANGLE WITH 2 SIDES = X & A DIAGONAL OF 6 METERS.
A^2+B^2=C^2
X^2+X^2=6^2
2X^2-36
X^2=36/2
X^2=18
X=4.2 ANS. FOR THE SIDES OF THE SQUARE.